# Memory-induced long-range order drag

### v1

@@ I Introduction
Long-range order (LRO) is typically associated with systems tuned to a critical point, where correlations span the entire size and collective dynamics emerge [1, 2, 3]. Recent work, however, has shown that LRO can also occur away from criticality when memory—in the form of slow, auxiliary degrees of freedom—interacts with fast, locally coupled variables [4]. This memory-induced LRO (MILRO) produces robust power-law correlations in the fast degrees of freedom, and has been observed in diverse contexts including spin-glass models [4], quantum spin chains [5], neuromorphic devices [6], models of neural activity [7], memcomputing machines [8, 9], and feedback-induced quantum phase transitions [10]. In these systems, slow memory variables act as an effective long-range interaction in time, inducing spatial LRO in the fast degrees of freedom over a wide range of parameters, and offering a mechanism for collective dynamics that does not require fine parameter tuning.
It is worth noticing that in all the above examples, the memory degrees of freedom share the same spatial region as the fast degrees of freedom. An open question is then whether such MILRO can propagate even beyond spatial regions of the system where the memory resides. Answering this question is not just an academic exercise. Rather, it is motivated by the fact that in many natural and engineered systems, functional units are arranged in modules or layers that interact through local couplings. A prototypical example is the brain architecture, where the neocortex is organized into distinct layers with structured feedforward and feedback connectivity, enabling hierarchical and laminar information processing [11, 12, 13, 14]. If a subset of these layers contains memory-rich dynamics, hence possibly supporting MILRO, can their order “drag” downstream into memory-free layers, allowing the entire stack to sustain collective, scale-free activity? Furthermore, could this effect explain the experimentally observed scale-free dynamics in the firing of neurons [15]?
@@ II.1 Model dynamics
Above, each 2D layer $k\geq 0$ consists of the standard, nearest-neighbor spin-glass interactions $J_{ij}^{(k)}=\pm J$, uniformly randomized in each layer. In the base layer $k=0$ only do we introduce an additional degree of freedom (DOF) $x_{ij}$ which grows or decays based on the state of its two accompanying spins, $s_{i}^{(0)}$ and $s_{j}^{(0)}$. In all other layers $k\geq 1$, there is an additional unidirectional ferromagnetic/anti-ferromagnetic interaction $J_{i}^{\perp(k)}=\pm J^{\perp}$ (again, assigned randomly) which couples $s_{i}^{(k)}$ to $s_{i}^{(k-1)}$. We clamp these dynamics so that $s_{i}^{(k)}\in[-1,1]$ and $x_{ij}\in[0,1]$. As in Ref. [4] and unless otherwise specified, we fix $\delta=3/4$, $g=2$ and $J=1$ for the remainder of this work. A schematic representing this model can be found in Fig. 1.
@@ II.2 MILRO Drag
where $\overline{f}$ denotes the average of some quantity $f$ over the interval $t\in[l/\gamma,(l+1)/\gamma)$. As this process is iterated further, additional, strong couplings begin to emerge between spins further and further from each other in the lattice. This induces an effective long-range interaction which emerges due to the presence of memory alongside the explicitly local spin-glass interactions [4].
@@ III Numerical verification of the MILRO drag
We now demonstrate the existence of MILRO drag numerically using avalanche statistics, supporting the previous theoretical analysis. Avalanches are defined as spatiotemporal bursts of spin flips separated by quiescent intervals; the size $s$ is the total number of flips in a burst. More numerical details are reported in Appendices B and C.
@@ III.1 Two-layer test and finite-size scaling
should hold [1]. Using the $\tau_{k}$ extracted from Fig. 2, we do indeed obtain excellent collapses of $s^{\tau_{k}}P_{k}(s)$ versus $s/L^{D}$ with $D=2$ for both $k=0$ and $k=1$ (insets of Fig. 2). This shows that the downstream, memory-free layer exhibits intra-layer scale-free avalanches with an area-controlled cutoff, which is a direct evidence of LRO drag.
Figure 3: Depth-resolved phase structure and representative avalanche statistics. (a) Avalanche size distributions $P(s)$ for $\{\gamma,J^{\perp}\}=\{0.045,5.0\},\{0.045,3.1\},\{2.2,3.1\}$ (stars correspond to locations in the phase diagram in panel b). Dashed lines indicate power-law guides with exponents $\gamma\simeq 1.90$ and $1.92$, respectively. (b) Phase diagrams in the $\{\gamma,J^{\perp}\}$ plane at depths $k=0,1,2,5,10$. Colors denote long-range order, crossover, and short-range correlations. Stars mark the parameter points used in (a). All points in the parameter space are simulated over 25 instances for $T=200$ simulation time (a.u.) (with initial transient dynamics removed) on a $64\times 64$ lattice.
@@ III.2 Phase diagram
Lastly, we must emphasize that for $J^{\perp}/J>4$ the inter-layer term always dominates and downstream layers largely follow the base layer’s dynamics, causing inter-layer synchronizations. This is precisely why the LRO region saturates in size after a few layers (see the similarity between Layer 5 and 10 in Fig. 3(b)). This is also why we expect the phase diagram not to change much if we were to simulate even deeper layers. On the contrary, for $J^{\perp}/J<2$ the system cannot overcome typical in-plane basins of attraction and layers quickly freeze. See Appendix A for a more rigorous analytical understanding of this point, and in Appendix D, we demonstrate the phase structures when we consider correlations within the entire three-dimensional stack instead of intra-layer only.
@@ III.3 Transient build-up of order
Figure 4: Transient build-up of long-range order across layers. Time evolution of the in-plane correlation length $\xi_{k}(t)$ for the driving layer ($k=0$) and ten downstream layers (k=110k=1-10) at $\{\gamma,J^{\perp}\}=\{0.045,5.0\}$ on a $64\times 64$ lattice with $J=1$, simulated over 100 instances for 4096 time steps, where each time step $\Delta t=0.017$. All layers exhibit a rapid rise over $\sim 10^{2}$ steps followed by a plateau. Mild overshoots and subsequent relaxation reflect the exclusion of system-wide events from $\xi$.
@@ IV Conclusion
These results suggest practical principles for multilayer neuromorphic systems: a memory-rich “source” layer can broadcast coherent, scale-free dynamics to passive layers without fine tuning. This mechanism further provides a concrete hypothesis for inter-laminar correlations in the brain cortex: slow adaptive processes, confined to specific layers, can sustain depth-wide coherence through local pathways. Based on these results, we expect this MILRO drag effect could be not just a mechanism of theoretical interest but also a design tool for layered, adaptive matter and computation.
Data availability—The code for reproducing the results in this paper can be found at https://github.com/chessonsipling/LRO_drag-main.
Acknowledgements.—We thank Kaushik Madapati for useful discussions. This work was funded by the National Science Foundation via grant No. ECCS-2229880. M.D. also acknowledges funding by the Alexander von Humboldt Stiftung through the Humboldt Research Award.
@@
Appendix A Analytical Derivation of Average Spin Flipping Time
@@
$=$ $\frac{1}{\Delta t}\Bigg(\frac{2\alpha}{4\Delta t}(\Delta t)^{4}+\bigg(\frac{2\beta}{3\Delta t}-\frac{\alpha}{3}\bigg)(\Delta t)^{3}$
$+\bigg(\frac{2x_{ij}(0)}{2\Delta t}-\frac{\beta}{2}\bigg)(\Delta t)^{2}-x_{ij}(0)(\Delta t)\Bigg),$
@@
Figure 5: The flipping time $\Delta t$ over which a single spin $s_{i}^{(k)}$ flips, in both a) the base layer $k=0$ and b) other layers $k\geq 1$. This timescale varies depending on the nature of the relevant spin-glass interactions. a) When $J_{4}=0$, $\Delta t=\sqrt{3/g\gamma\delta^{*}}$. When $J_{4}<0$, the flipping time decays as $\Delta t\sim 1/\gamma$, until $\gamma\sim O(1)$. When $J_{4}>0$, the flipping time remains constant at values near $1$ for $J_{4}=2J$ and $1/2$ for $J_{4}=4J$, until $\gamma\sim O(1)$. b) $\Delta t$ decays monotonically when both $J_{4}>0$ and $J_{1}^{\perp}>0$. Divergences appear at $J^{\perp}=2J$ and $J^{\perp}=4J$ when the inter-layer and intra-layer spin-glass interactions are in opposition, and in these cases, flips can only occur within a particular range of $J^{\perp}$. When $J_{4}<0$ and $J_{1}^{\perp}<0$, no flip will ever occur, so these curves cannot be plotted. Parameters chosen: $g=2$, $\delta^{*}\equiv(\delta-1/2)=1/4$, $J=1$.
Figure 6: The average time $\Delta t$ over which a single spin $s_{i}^{(k)}$ flips, in both a) the base layer $k=0$ and b) other layers $k\geq 1$. In both cases, all $\langle\Delta t\rangle$ shown in Fig. 5 are averaged according to the relative likelihood of occurrence, when both $J_{ij}^{(0)}s_{i}^{(0)}$ and $J_{i}^{\perp(k)}s_{i}^{(k-1)}$ are taken to be drawn from uniform distributions $\{-J,J\}$ and $\{-J^{\perp},J^{\perp}\}$, respectively. a) The curve $\langle\Delta t\rangle$ is fitted to a polynomial decay for $\gamma\in[10^{-2},10^{-1}]$, yielding $\Delta t\sim\gamma^{-0.96}$, and $\gamma\in[10^{1},10^{2}]$, yielding $\Delta t\sim\gamma^{-0.53}$. If this plot were to be extended, $\langle\Delta t\rangle$ would tend further towards $\Delta t\sim\gamma^{-1}$ for lower $\gamma$ and $\Delta t\sim\gamma^{-1/2}$ for higher $\gamma$. Parameters chosen: $g=2$, $\delta^{*}\equiv(\delta-1/2)=1/4$, $J=1$.
@@
Appendix B Numerical details
Appendix C Choosing the temporal proximity window
Figure 9: The optimal temporal proximity windows $\Delta t_{w}$ for different parameter configurations. Each data point is obtained by simulating the dynamics of 1 instance on a $64\times 64$ lattice over 8000 time steps ($T=8000\Delta t$, $\Delta t=0.048\gamma^{1/3}$), extracting the avalanches using 20 different values of $\Delta t_{w}$ and selecting the one that maximizes the correlation length. We can loosely see the phase structures presented in Fig. 3. We also note a line of large $\Delta t_{w}$ at $J^{\perp}=4$, corresponding to the diverging average spin flip time derived in Appendix A.
Appendix D Three–dimensional avalanche structures
The analysis in the main text focused on in–plane avalanches, i.e., clusters built from nearest–neighbor connectivity within a single layer. To probe inter-layer correlations directly, we extend the definition of an avalanche to three dimensions: two flip events are considered connected if they occur on nearest–neighbor sites in the full $(x,y,k)$ lattice and within the temporal window $\Delta t_{w}$ (cf. Appendix B). Therefore clusters may percolate both laterally and across layers.

### v2

@@ I Introduction
Long-range order (LRO) is typically associated with systems tuned to a critical point, where correlations span the entire system and collective dynamics emerge [26, 20, 12]. Recent work, however, has shown that LRO can also occur away from criticality when memory—in the form of slow, auxiliary degrees of freedom—interacts with fast, locally coupled variables [22]. This memory-induced LRO (MILRO) produces robust power-law correlations in the fast degrees of freedom, and has been observed in diverse contexts including spin-glass models [22], quantum spin chains [28], neuromorphic devices [29], models of neural activity [25], memcomputing machines [7, 23], and feedback-induced quantum phase transitions [13]. In these systems, slow memory variables act as an effective long-range interaction in time, inducing spatial LRO in the fast degrees of freedom over a wide range of parameters, and offering a mechanism for collective dynamics that does not require fine parameter tuning.
It is worth noticing that in all the above examples, the memory degrees of freedom share the same spatial region as the fast degrees of freedom. An open question is then whether such MILRO can propagate even beyond spatial regions of the system where the memory resides. Answering this question is not just an academic exercise. Rather, it is motivated by the fact that in many natural and engineered systems, functional units are arranged in modules or layers that interact through local couplings. A prototypical example is the brain architecture, where the neocortex is organized into distinct layers with structured feedforward and feedback connectivity, enabling hierarchical and laminar information processing [11, 8, 18, 9]. If a subset of these layers contains memory-rich dynamics, hence possibly supporting MILRO, can their order “drag” downstream into memory-free layers, allowing the entire stack to sustain collective, scale-free activity? [^1] Furthermore, could this effect explain the experimentally observed scale-free dynamics in the firing of neurons [3]?
@@ II Analytical understanding of the MILRO drag
We note that this model is an extension of the one used in Ref. [22], which considers only a single layer of spins with memories. That study acts as an essential background for the current work.
@@ II.1 Model dynamics
Here, $s_{i}^{(k)}$ denotes the value of the spin variable at layer $k$, planar coordinate $i$. Each 2D layer $k\geq 0$ consists of the standard, nearest-neighbor spin-glass interactions $J_{ij}^{(k)}=\pm J$, uniformly randomized in each layer; $\sum_{\langle ij\rangle}$ denotes summation over 2D nearest-neighbors. In the base layer $k=0$ only do we introduce an additional degree of freedom (DOF) $x_{ij}$ which grows or decays based on the state of its two accompanying spins, $s_{i}^{(0)}$ and $s_{j}^{(0)}$. In all other layers $k\geq 1$, there is an additional unidirectional ferromagnetic/anti-ferromagnetic interaction $J_{i}^{\perp(k)}=\pm J^{\perp}$ (again, assigned randomly) which couples $s_{i}^{(k)}$ to $s_{i}^{(k-1)}$. We clamp these dynamics so that $s_{i}^{(k)}\in[-1,1]$ and $x_{ij}\in[0,1]$. As in Ref. [22] and unless otherwise specified, we fix $\delta=3/4$, $g=2$ and $J=1$ for the remainder of this work; $\gamma$ and $J^{\perp}$ are tunable parameters that we will study in the following sections. All summations are provided explicitly in these dynamical equations and following analytics; Einstein notation is used nowhere. A schematic representing this model can be found in Fig. 1.
@@ II.2 MILRO Drag
where $\overline{f}$ denotes the average of some quantity $f$ over the interval $t\in[l/\gamma,(l+1)/\gamma)$. As this process is iterated further, additional, strong couplings begin to emerge between spins farther and farther from each other in the lattice. This induces an effective long-range interaction which emerges due to the presence of memory alongside the explicitly local spin-glass interactions [22].
@@ III Numerical verification of the MILRO drag
We now demonstrate the existence of MILRO drag numerically using avalanche statistics, supporting the previous theoretical analysis. Avalanches are defined as spatiotemporal bursts of spin flips separated by quiescent intervals; the size $s$ is the total number of flips in a burst. More numerical details are reported in Appendices A and B.
@@ III.1 Two-layer test and finite-size scaling
should hold [26]. Using the $\tau_{k}$ extracted from Fig. 2, we do indeed obtain excellent collapses of $s^{\tau_{k}}P_{k}(s)$ versus $s/L^{D}$ with $D=2$ for both $k=0$ and $k=1$ (insets of Fig. 2). This shows that the downstream, memory-free layer exhibits intra-layer scale-free avalanches with an area-controlled cutoff, which is a direct evidence of LRO drag. Furthermore, this downstream scale-invariance emerges even in the presence of quenched disorder, in the form of randomized inter-layer couplings $J_{i}^{\perp(k)}=\pm J^{\perp}$, and heterogeneous memory timescales in the base layer as shown explicitly in Appendix C.
@@ III.2 Phase diagram
Figure 3: Depth-resolved phase structure and representative avalanche statistics. (a) Avalanche size distributions $P(s)$ for $\{\gamma,J^{\perp}\}=\{0.045,5.0\},\{0.045,3.1\},\{2.2,3.1\}$ (stars correspond to locations in the phase diagram in panel (b). Dashed lines indicate power-law guides with exponents $\gamma\simeq 1.90$ and $1.92$, respectively. (b) Phase diagrams in the $\{\gamma,J^{\perp}\}$ plane at depths $k=0,1,2,5,10$. Colors denote long-range order, crossover, and short-range correlations. Stars mark the parameter points used in (a). All points in the parameter space are simulated over 25 instances for $T=200$ simulation time (a.u.) (with initial transient dynamics removed) on a $64\times 64$ lattice.
@@
Lastly, we must emphasize that for $J^{\perp}/J>4$ the inter-layer term always dominates and downstream layers largely follow the base layer’s dynamics, causing inter-layer synchronizations. This is precisely why the LRO region saturates in size after a few layers (see the similarity between Layer 5 and 10 in Fig. 3(b)). This is also why we expect the phase diagram not to change much if we were to simulate even deeper layers. On the contrary, for $J^{\perp}/J<2$ the system cannot overcome typical in-plane basins of attraction and layers quickly freeze. See Appendix D for a more rigorous analytical understanding of this point, and in Appendix E, we demonstrate the phase structures when we consider correlations within the entire three-dimensional stack instead of intra-layer only. It is also worth noticing that the transfer of LRO does not arise trivially from the forcing of spins in lower layers to faithfully track the dynamics of their upper counterparts. We explicitly show this in Appendix F.
@@ III.3 Transient build-up of order
Figure 4: Transient build-up of long-range order across layers. Time evolution of the in-plane correlation length $\xi_{k}(t)$ for the driving layer ($k=0$) and five downstream layers ($k=2,4,6,8,10$) at $\{\gamma,J^{\perp}\}=\{0.045,5.0\}$ on a $64\times 64$ lattice with $J=1$, simulated over 100 instances for 4096 time steps, where each time step $\Delta t=0.017$. All layers exhibit a rapid rise over $\sim 10^{2}$ steps followed by a plateau. Mild overshoots and subsequent relaxation reflect the exclusion of system-wide events from $\xi$. Note that this transient timescale is $\sim 1/\gamma$, as expected, since $1/\gamma$ acts as an approximate timescale for the memory dynamics.
@@ IV Conclusion
These results suggest practical principles for multilayer neuromorphic systems: a memory-rich “source” layer can broadcast coherent, scale-free dynamics to passive layers without fine tuning. This mechanism further provides a concrete hypothesis for inter-laminar correlations in the brain cortex, readily testable with existing volumetric neural imaging techniques [21, 6, 10]: slow adaptive processes, confined to specific layers, can sustain depth-wide coherence through local pathways. Based on these results, we expect this MILRO drag effect could be not just a mechanism of theoretical interest but also a design tool for layered, adaptive matter and computation.
Acknowledgments
We thank Kaushik Madapati for useful discussions. This work was funded by the National Science Foundation via grant No. ECCS-2229880. M.D. also acknowledges funding by the Alexander von Humboldt Stiftung through the Humboldt Research Award.
Author contributions
@@ Acknowledgments
Appendix A Numerical details
Appendix B Choosing the temporal proximity window
Figure 7: The optimal temporal proximity windows $\Delta t_{w}$ for different parameter configurations. Each data point is obtained by simulating the dynamics of 1 instance on a $64\times 64$ lattice over 8000 time steps ($T=8000\Delta t$, $\Delta t=0.048\gamma^{1/3}$), extracting the avalanches using 20 different values of $\Delta t_{w}$ and selecting the one that maximizes the correlation length. We can loosely see the phase structures presented in Fig. 3. We also note a line of large $\Delta t_{w}$ at $J^{\perp}=4$, corresponding to the diverging average spin flip time that will be derived in Appendix D.
Appendix C Heterogeneous Memory Timescales
To test how the emergence of LRO drag is influenced by the presence of spatially heterogeneity in the base layer, we perform an additional set of simulations in which the memory parameter $\gamma$ in Eqn. 1 now takes a unique value, $\gamma_{ij}$, for each edge in the base layer, rather than this constant being fixed at each edge. These parameters $\gamma_{ij}$ are independently sampled from a Gaussian distribution with mean $\mu_{\gamma}$ and standard deviation $\sigma_{\gamma}$. Parameters $\gamma_{ij}$ are subsequently bounded from below by $10^{-3}$ to assure that these frequencies remain positive. Note that each memory variable still evolves according to a dynamical equation with the same functional form given in Eqn. 1, but $\gamma$ now takes a distinct value at each edge.
Fig. 8 shows the resulting avalanche distributions in layers 0 and 1 for Gaussians with mean $\mu_{\gamma}=0.05$ and two different standard deviations $\sigma_{\gamma}=0.005$ and $\sigma_{\gamma}=0.02$. In both cases, scale-free avalanche distributions in layer 0 are transferred to layer 1 despite the fact that individual memories evolve over different dynamical timescales. This further suggests that no fine-tuning is present in our model, and that no precise functional form of the memories is required to generate or transfer LRO. So long as the majority of memories evolve much more slowly than their accompanying spins (which is guaranteed so long as $\mu_{\gamma}$ is not too large), LRO will be produced that can then be dragged to subsequent layers.
Figure 8: Avalanche distributions with randomized memory parameter $\gamma_{ij}$ in the base layer, where each $\gamma_{ij}$ is independently drawn from a Gaussian with mean $\mu_{\gamma}=0.05$ and standard deviation $\sigma_{\gamma}$. Avalanches are extracted at both $\sigma_{\gamma}=0.005$ (a) and $\sigma_{\gamma}=0.02$ (b). In each case, scale-free avalanches persist in both the base layer 0 and layer 1, despite the heterogeneity of the memory timescales. Avalanches are extracted over 1000 instances, each simulated for $T=200$ simulation time (a.u.), on a $64\times 64$ lattice with $\Delta_{tw}/dt=125$, and $J^{\perp}=3.0$.
Appendix D Analytical Derivation of Average Spin Flipping Time
@@
$=$ $\frac{1}{\Delta t}\Bigg(\frac{2\alpha}{4\Delta t}(\Delta t)^{4}+\bigg(\frac{2\beta}{3\Delta t}-\frac{\alpha}{3}\bigg)(\Delta t)^{3}+\bigg(\frac{2x_{ij}(0)}{2\Delta t}-\frac{\beta}{2}\bigg)(\Delta t)^{2}-x_{ij}(0)(\Delta t)\Bigg),$
@@
Figure 9: The flipping time $\Delta t$ over which a single spin $s_{i}^{(k)}$ flips, in both (a) the base layer $k=0$ and (b) other layers $k\geq 1$. This timescale varies depending on the nature of the relevant spin-glass interactions. (a) When $J_{4}=0$, $\Delta t=\sqrt{3/g\gamma\delta^{*}}$. When $J_{4}<0$, the flipping time decays as $\Delta t\sim 1/\gamma$, until $\gamma\sim O(1)$. When $J_{4}>0$, the flipping time remains constant at values near $1$ for $J_{4}=2J$ and $1/2$ for $J_{4}=4J$, until $\gamma\sim O(1)$. (b) $\Delta t$ decays monotonically when both $J_{4}>0$ and $J_{1}^{\perp}>0$. Divergences appear at $J^{\perp}=2J$ and $J^{\perp}=4J$ when the inter-layer and intra-layer spin-glass interactions are in opposition, and in these cases, flips can only occur within a particular range of $J^{\perp}$. When $J_{4}<0$ and $J_{1}^{\perp}<0$, no flip will ever occur, so these curves cannot be plotted. Parameters chosen: $g=2$, $\delta^{*}\equiv(\delta-1/2)=1/4$, $J=1$.
Figure 10: The average time $\Delta t$ over which a single spin $s_{i}^{(k)}$ flips, in both (a) the base layer $k=0$ and (b) other layers $k\geq 1$. In both cases, all $\langle\Delta t\rangle$ shown in Fig. 9 are averaged according to the relative likelihood of occurrence, when both $J_{ij}^{(0)}s_{i}^{(0)}$ and $J_{i}^{\perp(k)}s_{i}^{(k-1)}$ are taken to be drawn from uniform distributions $\{-J,J\}$ and $\{-J^{\perp},J^{\perp}\}$, respectively. (a) The curve $\langle\Delta t\rangle$ is fitted to a polynomial decay for $\gamma\in[10^{-2},10^{-1}]$, yielding $\Delta t\sim\gamma^{-0.96}$, and $\gamma\in[10^{1},10^{2}]$, yielding $\Delta t\sim\gamma^{-0.53}$. If this plot were to be extended, $\langle\Delta t\rangle$ would tend further towards $\Delta t\sim\gamma^{-1}$ for lower $\gamma$ and $\Delta t\sim\gamma^{-1/2}$ for higher $\gamma$. Parameters chosen: $g=2$, $\delta^{*}\equiv(\delta-1/2)=1/4$, $J=1$.
@@
Appendix E Three–dimensional avalanche structures
The analysis in the main text focused on in–plane avalanches, i.e., clusters built from nearest–neighbor connectivity within a single layer. To probe inter-layer correlations directly, we extend the definition of an avalanche to three dimensions: two flip events are considered connected if they occur on nearest–neighbor sites in the full $(x,y,k)$ lattice and within the temporal window $\Delta t_{w}$ (cf. Appendix A). Therefore clusters may percolate both laterally and across layers.
@@ Appendix E Three–dimensional avalanche structures
Appendix F Absence of Forcing Between Subsequent Layers
From our previous results, it is still possible that the transfer of LRO arises trivially from the forcing of spins in lower layers to precisely track the dynamics of their upper counterparts, regardless of the dynamics of these lower layers. To dispel this possibility, we calculate the correlation between flipping spins in adjacent layers $k$ and $k+1$ explicitly, defined as follows:
$\tilde{G}^{(k)}(\tau)=\Big\langle\text{sgn}\Big(J_{i}^{\perp(k)}\Big)\dot{s_{i}}^{(k)}(t)\dot{s_{i}}^{(k+1)}(t+\tau)\Big\rangle\,.$ (16)
Above, the average is first taken over space, at all indices $i$, and then over time, at all instances $t$ when the upper layer’s spin is actively flipping (i.e., its derivative is non-zero). This function explicitly depends on $\tau$, a measure of the time delay between spin flips in adjacent layers. The factor of $\text{sgn}(J^{\perp})$, where $\text{sgn}(\cdot)$ returns either +1+1 or 1-1 based on the sign of its argument, is incorporated since trivial forcing corresponds to spins $s_{i}^{(k)}$ and $s_{i}^{(k+1)}$ flipping oppositely if their interaction is anti-ferromagnetic.
In Fig. 12, Eqn. 16 is plotted as a function of time delay $\tau$ when $J^{\perp}=3.0$ and $8.0$. When $J^{\perp}$ is not too large, $\tilde{G}$ peaks far below unity after a short time delay before dropping to approximately zero for longer delays. This suggests that these lower-layer spins do not trivially track the dynamics of the upper spins, although the directional flow of information in time can still be resolved. On the other hand, when $J^{\perp}=8.0$, this peak occurs nearly instantaneously and takes a much higher value. The insets in Fig. 12 showing individual spin dynamics makes the following especially clear: so long as $J^{\perp}$ is not so large as to dominate over intra-layer interactions, lower layers do not trivially follow the dynamics of their upper neighbors. Nevertheless, LRO can still be dragged through the layers.
Figure 12: The correlation function $\tilde{G}^{(1)}$ as a function of $\tau$ for (a) $J^{\perp}=3.0$ and (b) $J^{\perp}=8.0$. In both cases, the memory frequency $\gamma=0.05$. A random sample of 5 spin trajectories are shown in the inset, with spins in the same lattice position $i$ in adjacent layers plotted in the same color. (a) $\tilde{G}$ peaks far below unity before dropping to near zero, suggesting spins in layer 2 are not being trivially forced to track the dynamics of layer 1 spins. In the inset, several layer 1 spin flips do not induce layer 2 flips at all. (b) $\tilde{G}$ peaks much closer to unity. All spin trajectories in layer 2 appear to trivially follow those in layer 1.